Project

General

Profile

03062014 Minutes

Attending: Micheal, Mat, Gavin, Dom, Evan

Michael - updated the nearline wiki
Pavan, Siva - nearline checklist is in works

Diblock 7 going cold today.

12 diblock testing yesterday,

water leak last night

hits/spill was shown at outfitting. Bill Miller - is notified. Might be air handling?

NDOS ran out-of-sync due to TDUControl issue - Fixed this morning.

We should add the watchdog page to the DAQ/Computing form.